# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v1

@@ 5.1 Stability and Attainability of Sobolev Levels
Thus, we prove the inequality and $\mathcal{G}_{3,1}[H_{3,1}]=(\frac{1}{4},\infty).$ By [25, Theorem 1.7], we obtain $H_{3,1}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
@@
(i) By [20, Theorem 1.14] and [23, Theorem 1.7], we know that $H_{n,k}(\lambda)<S_{n,k}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $n\geq 4k$. By arguments completely analogous to those used in the preceding propositions in high dimensions, we obtain the desired conclusion here, and therefore omit the details.
(ii) By [20, Theorem 1.14] and [23, Theorem 1.7], there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $H_{n,k}(\lambda)<S_{n,k}$ when $\lambda>\lambda_{k}^{\mathrm{conf}}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$. The remainder of the proof follows by an entirely analogous argument, combined with Lemma 5.6 (ii) and Proposition 4.11, and we therefore omit the details.
(iii) When $n=2k+1$, by [24, Theorem 1.6] or [23, Theorem 1.2], we obtain $H_{n,k}(\lambda_{0,k}^{\mathrm{conf}})=S_{n,k}.$ The rest of the proof is completely analogous to (ii), and we omit it. $\Box$

### v2

@@ 5.1 Stability and Attainability of Sobolev Levels
Thus, we prove the inequality and $\mathcal{G}_{3,1}[H_{3,1}]=(\frac{1}{4},\infty).$ By [30, Theorem 1.7], we obtain $H_{3,1}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
@@
(i) By [25, Theorem 1.14] and [28, Theorem 1.7], we know that $H_{n,k}(\lambda)<S_{n,k}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $n\geq 4k$. By arguments completely analogous to those used in the preceding propositions in high dimensions, we obtain the desired conclusion here, and therefore omit the details.
(ii) By [25, Theorem 1.14] and [28, Theorem 1.7], there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $H_{n,k}(\lambda)<S_{n,k}$ when $\lambda>\lambda_{k}^{\mathrm{conf}}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$. The remainder of the proof follows by an entirely analogous argument, combined with Lemma 5.5 (ii) and Proposition 4.11, and we therefore omit the details.
(iii) When $n=2k+1$, by [29, Theorem 1.6] or [28, Theorem 1.2], we obtain $H_{n,k}(\lambda_{0,k}^{\mathrm{conf}})=S_{n,k}.$ The rest of the proof is completely analogous to (ii), and we omit it. $\Box$
Conflicts of interest: The authors declare that they have no conflicts of interest regarding this work.
Data availability: This paper has no associated data.
Acknowledgements: H. Chen is supported by NSFC, no. 12361043.
R. Chen is supported by China Scholarship Council, Liujinxuan [2025] no. 37.